Bwindi Impenetrable Forest (01.01.2014 to 03.31.2014)

You are here

PDF versionPDF version
Year: 
2014
Quarter: 
January - March
Communications activities: 

Badru Mugerwa attended a student's conference on conservation science.

Unusual Events at the TEAM Site: 

A feral cat was recorded on a camera trap at the TEAM site. Initial speculation was that the recorded cat is an African wildcat (Felis silvestris). Expert consultation confirmed the species as a domestic-feral cat.

New Species: 

No new species recorded at TEAM site.

Protocol Activities: 

Only vegetation and climate monitoring protocols were implemented during this quater. Two vegetation plots were remeasured in January. Data was uploaded onto the TEAM portal.

Climate data and  bi-weekly maintenance logs were also uploaded onto the portal.

Protocol Problems: 

Vegetation monitoring protocol

No protocol problems were encounted with the vegetation protocol.

Climate monitoring protocol

A Pyronometer was recalibrated by Scientific Campbell, and shipped back to Uganda. A severe corrosion damage beyond repair was found on the circuits for both HMP45C sensors. The two sensors were declared non-repairable by the Scientific Campbell technicians.  It was advised that they get replaced. The HMP45Cs are obsolete and are not manufactured any more, hence will be replaced by the HC2-S3 sensors. The station is currently running on only one humidity/temperature sensor. The two HC2-S3 sensors are not yet installed due to lack of funds.  A gradual decrease in battery voltage has also been observed from the monthly climate data. The cause is not yet known. Battery voltages have shown a similar trend during past cloudy/rainy seasons. I anticipate the low battery voltage could be due to the cloudy conditions of the wet/rainy season.  The two  humidity/temperature sensor will be installed, together with a  second solar panel soonest in the next financial year (starting July 2014). Climate data for this quarter was submitted to the TEAM portal. Bi-weekly maintenance of the climate station has been carried out in accordance with the protocol. The maintenance logs were also uploaded onto the portal.

Terrestrial vertebrate monitoring protocol

No terrestrial vertebrate monitoring activity was done in this quater. However, I suggest that at least ten camera traps are replaced to avoid future camera failure and eventual data loss.  It is not possible to have the camera traps replaced by the time of the survey due to unavailability of funds.

 

Schedule Problems: 

No schedule problems ecncountered during protocol implementation. All activities were implemented according to schedule.

Logistical Problems: 

Replacing the broken humidity/temperature sensors and atleast ten camera traps has not been possible due to luck of funds in the current financial year's budget. These costs will be included in the next year's fimacial year starting July 2014 to allow these relacements. The  HC2-S3 sensors cost $644 including cable and taxes per piece. The HC500 HyperFire Semi-Covert IR camera traps cost $ 986.9 per piece (including accessories; security enclosure, cable locks, SD cards and taxes) .

Suggestions: 

This is not a suggestion, but rather sharing progress on the edge effect analyses at TEAM sites which Doug, Melissa and I are working on. We are progressing well. Edge data for BIF and UDZ has been compiled. Next to contact are site managers of Pasoh, Manu and Volcan Barva TEAM sites. So far all is well. I will share the paper outline once we have all the data in place.